OpenStudy (yttrium):

A branch falls from the top of a 95.0m tall redwood tree, starting from rest. Use work-kinetic energy theorem to determine the speed of the branch when it reaches the ground.

OpenStudy (yttrium):

@Hero @Callisto

OpenStudy (yttrium):

@ash2326 @chemENGINEER

OpenStudy (yttrium):

@thomaster

ganeshie8 (ganeshie8):

work done = change in KE apply it ?

OpenStudy (yttrium):

Then it should be W = mV^2/2 - mVo^2/2

OpenStudy (math&ing001):

Your v0=0

ganeshie8 (ganeshie8):

Yup ! plug the values and W = F.s

OpenStudy (yttrium):

I have no initial velocity, hence. W= mV^2/2, is that right?

ganeshie8 (ganeshie8):

Correct

OpenStudy (yttrium):

Then I will have mad = mV^2/2 ad = V^2/2, where a=g, right? Hence V = sqrt(2gd) V = sqrt(2(9.8)(95)) Resulting to V = 43.15m/s Did I've done it correctly?

OpenStudy (math&ing001):

Yes

ganeshie8 (ganeshie8):

yup ! thats correct

ganeshie8 (ganeshie8):

you can verify it using the kinematics if u want, until u get hang of conservation of energy concept

ganeshie8 (ganeshie8):

v^2 - u^2 = 2ad put u = 0 you wil get the same equation v = sqrt(2gd)

OpenStudy (yttrium):

Yeah @ganeshie8 . Thanks >:)). I still have one more, can you guide me?

ganeshie8 (ganeshie8):

il try :)

OpenStudy (yttrium):

Shot put competitions use metal balls with a mass of 16lb. A competitor throws the shut at an angle of 43.3 deg and releases it from a height of 1.82m above where it lands, and lands a horizontal distance of 17.7m from the point of release. What is the kinetic energy of the shot as it leaves the thrower's hand?

ganeshie8 (ganeshie8):